OpenStudy (anonymous):

Which graph shows the quadratic function y = 3x2 − 12x + 10? A B C D

OpenStudy (anonymous):

heres the graphs

OpenStudy (anonymous):

@Callisto

OpenStudy (callisto):

1. Check delta delta = (-12)^2 - 4(3)(10) = 24 >0 => 2 intercepts with x-axis 2. find vertex 3x2 − 12x + 10 = 3(x^2 -4x) +10 = 3(x-2)^2 -12 +10 = 3(x-2)^2 -2 So, vertex is at (2, -2) Can you get the answer now?

OpenStudy (anonymous):

I'm confused as to would this be the 2nd graph or the 4th @Callisto

OpenStudy (anonymous):

I know this is like really old but ...

OpenStudy (callisto):

What do you think?

OpenStudy (anonymous):

My Hearts Telling me D But My Mind Is Telling Me B

OpenStudy (callisto):

Why is your heart telling you it's D?

OpenStudy (anonymous):

Because It Looks Like It's Set in the right spot , But then again my heart has been broken so many times Thats Its not sure....

OpenStudy (callisto):

And why is your mind telling you it's B?

OpenStudy (anonymous):

My Mind Is Telling Me It's B Because , It Looks Like It Goes Up 2 then slide to -2 , Looks More like a rise over run

OpenStudy (callisto):

Do you know what does "x-intercept" mean?

OpenStudy (anonymous):

Like x would be 2

OpenStudy (callisto):

What if I say there are two x-intercepts, what does that mean?

OpenStudy (anonymous):

O_O I think the dots would be on the same line 

OpenStudy (anonymous):

I'm starting to think about it and I believe it is d

OpenStudy (callisto):

 Number of intercepts = how many times you "cut" or "touch" the x-axis.